OpenStudy (okdutchman7):

Please help. WILL MEDAL! Use the letter d and the number 100 as required to write an expression that tells the number of 3-cent stamps that can be purchased for d dollars. @jim_thompson5910 @sleepyjess @triciaal @hartnn

OpenStudy (mathstudent55):

If you have d dollars, how many cents do you have?

OpenStudy (okdutchman7):

100d cents

OpenStudy (mathstudent55):

Correct. How many 3-cent stamps can you buy with 100d cents?

OpenStudy (okdutchman7):

I'm not sure.

OpenStudy (mathstudent55):

OK. Let's start with examples that will lead to an answer. Let's say you have 3 cents. How many 3-cent stamps can you buy with 3 cents?

OpenStudy (okdutchman7):

1

OpenStudy (mathstudent55):

Good. Now think of what operation do you do to 3 cents (you have) and 3 cents (cost of stamp) that gives you an answer of 1? What operation do you do to 3 and 3 to get 1 as an answer?

OpenStudy (okdutchman7):

Division

OpenStudy (mathstudent55):

Excellent. Let's see if the same operation will work for other examples. With 6 cents, you can buy 2 3-cent stamps. 6/3 = 2 With 12 cents, you can buy 4 3-cent stamps. 12/3 = 4 As you can see, division works in all these examples.

OpenStudy (okdutchman7):

I see.

OpenStudy (okdutchman7):

Is it \[\frac{ 100d }{ 3 } stamps\]

OpenStudy (mathstudent55):

In every case above, we divided the amount of money we have by the cost of 1 stamp (3 cents in this case). You can expect that division will work for all examples. If you have 100d cents, how many 3-cent stamps can you buy? Divide the amount of money you have by the cost of 1 stamp. What do you get?

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Good job!

OpenStudy (okdutchman7):

Thank you for your help.

OpenStudy (mathstudent55):

You're welcome.